6)Two plots at Rothamsted Experimental Station were studied for production of wheat straw. For a random sample of years, the annual wheat straw production (in pounds) from one plot was as follows.

6.68 5.84 7.03 6.89 7.31 7.18
7.06 5.79 6.24 5.91 6.14

Use a calculator to verify that, for this plot, the sample variance is s2 ≈ 0.335.

Another random sample of years for a second plot gave the following annual wheat production (in pounds).

7.59 6.12 6.54 7.80 7.22 5.58 5.47 5.86

Use a calculator to verify that the sample variance for this plot is s2 ≈ 0.836.

Test the claim that there is a difference (either way) in the population variance of wheat straw production for these two plots. Use a 5% level of signifcance.




(b) Find the value of the sample F statistic. (Use 2 decimal places.)


What are the degrees of freedom?

dfN
dfD
Answer #1

b)

value of the sample F statistic. =(s2/s1)2 =(0.836/0.335)=2.50

degrees of freedom

dfN =8-1 =7

dfD =11-1 =10